# Guided Discrete Diffusion for Electronic Health Record Generation

### v1

@@ 6 Conclusion and Future Work
In this paper, we introduced a novel generative model for synthesizing realistic EHRs EHR-D3PM. Leveraging the latest advancements in discrete diffusion models, EHR-D3PM overcomes the challenges of GAN-based approaches and effectively generates high-quality tabular medical data. Compared with other diffusion-based approaches, EHR-D3PM enables high-quality conditional generation. Our experiment demonstrates that EHR-D3PM not only achieves state-of-the-art performance in fidelity, utility, and privacy metrics but also significantly improves downstream task performance through data augmentation. Further investigations of the vulnerability of diffusion-based generative models in EHR generation, particularly to Membership Inference Attacks (MIAs) (Shokri et al., 2017), is a promising future direction as well as providing formal privacy guarantees, e.g., by incorporating differential privacy, which is largely unexplored in diffusion-based models for discrete data.
@@
In the following Table 4, we present a concise summary of various diseases along with their corresponding International Classification of Diseases, Ninth Revision (ICD 9) codes. $\mathcal{D}_{1},\mathcal{D}_{2}$ represents the real dataset 1 and real dataset 2, respectively. This table includes common conditions such as Type II Diabetes (T2D), Chronic Kidney Disease (CKD), Chronic Obstructive Pulmonary Disease (COPD), Asthma, Hypertension and Osteoarthritis. Each disease is associated with specific ICD 9 codes that are used for clinical classification and diagnosis purposes. In this paper, we are interested in the diseases listed in Table 4.
@@
Disease ICD 9 Code MIMIC Dataset $\mathcal{D}_{1}$ Dataset $\mathcal{D}_{2}$
@@ A.1 Dataset Details
The MIMIC III dataset includes a patient population of 46,520. There are 651,047 positive codes within 64,314 hospital admission records (HADM IDs). We have implemented an 80/20 split for training and testing purposes. Specifically, this allocates 12,862 records for testing and the remaining 51,451 for training. The histograms in Fig. 2 indicate the density distribution of feature number per record. The dimension is $N=1042.$
@@
The first dataset, denoted by $\mathcal{D}_{1}$, includes a patient population of size 1,670,347. We split the whole dataset into 100K for validation, 2000K for testing and the rest 1,370, 347 for training. The number of codes per patient is relatively small, as indicated by the histogram of feature number per record in Fig. 4. The dimension is $N=993.$
@@
The second dataset, denoted by $\mathcal{D}_{2}$, includes a patient population of size 1,859,536. We split the whole dataset into 100K for validation, 2000K for testing and the rest 1,559,536 for training. The number of codes per patient is relatively large, as indicated by the histogram of feature number per record in Fig. 8. Although, dataset $\mathcal{D}_{2}$ has relatively denser feature, the prevalence of six chronic diseases we are interested in is pretty low. The dimension is $N=993.$
@@ A.2 Model Architecture Detail
The denoise model in this paper has a uniform architecture, illustrated in Fig. 6. The architecture we propose is tailored for tabular EHRs as it is non-sequential data. While the architecture proposed in multinomial diffusion Hoogeboom et al. (2021b) is designed for sequential data, where neighboring dimensions have semantic correlation. The tabular EHR datasets in our paper don’t have such property. Therefore, we propose a novel transformer-based model for tabular EHRs. One bottleneck of transformer models is that the computational complexity of the attention module is quadratic to the dimension of input data. We adopt an efficient block based on  Wang et al. (2020), whose attention operation has linear complexity with respect to the dimension of input data.
@@ A.3 Hyper-parameters
In the optimization phrase, we adopt adamW optimizer, and the weight decay in adamW is 1.e-5. The learning rate is 1e-4 and batch size is 256. The beta for expentialLR in learning rate schedule is 0.99. The number of training epochs is 100. It takes less than three hours to finish training this model on A6000 with 48G memory.
@@
The optimization parameters for both datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are also the same. In the optimization phrase, we adopt adamW optimizer. The learning rate is 1e-4 and batch size is 512. The weight decay in adamW is 1.e-5. The beta for expentialLR in learning rate schdule is 0.99. The number of training epochs is 40. It takes one and half day to train one model on A100 with 80G memory.
@@ A.4 Evaluation Metrics
$\widehat{\mathrm{MMD}}_{k_{\gamma}}(P,Q)=\frac{1}{n(n-1)}[\sum_{i\neq j}k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})+\sum_{i\neq j}k_{\gamma}(\mathbf{x}^{\prime}_{i},\mathbf{x}^{\prime}_{j})]-\frac{1}{n^{2}}\sum_{i,j}k_{\gamma}(\mathbf{x}_{i},\mathbf{x}^{\prime}_{j}).$
@@
$k_{\gamma}(\mathbf{x},\mathbf{x}^{\prime})=\mathrm{exp}(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2h^{2}_{\gamma}})$
@@ B.1 Additional experiments on unconditional generation
Fig. 7 provides additional comparison of marginal distribution matching on dataset $\mathcal{D}_{2}$. Since dataset $\mathcal{D}_{2}$ has relatively denser features, the performance of baselines in low prevalence regime is less severe than that of baselines on dataset $\mathcal{D}_{1}$. From the Spearman correlation in low prevalence regime, we can still see that our method significantly outperform baselines. Based on results in Fig. 1, Fig. 3 and Fig. 7, we consistently observe synthetic data by EHRDiff fails to capture the information in low prevalence regime. One reason we articulate is that the foundation of EHRDiff is designed for continuous distribution and cannot be readily applied to the generation of discrete data particularly on data of low prevalence regime. From Fig. 8, we can see that the histogram of feature number per record on synthetic data by our method provides a perfect matching to that of real data.
@@
We also apply our methods to downstream prediction tasks on dataset $\mathcal{D}_{2}$, where the prevalence of six chronic diseases is much lower. From Fig 5, we can see that the accuracy of our prediction is still close to the prediction of classifier models trained on real data, which is the classifier baseline. While other baselines have a much larger performance gap when compared with the ideal classifier. Particularly on rare diseases such as hypertension heart, the classifier trained on synthetic data by our model has 8% absolute improvement in AUPRC and AUROC over the strongest baseline on both $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. From the confidence intervals provided in Table 7, 6, 8 and 9, we confirm that such improvement over the baselines are statistically significant.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Spearm corr = 0.96 (b) Spearm corr = 0.99 (c) Spearm corr = 0.95 (d) Spearm corr = 0.99
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Spearm corr = 0.95 (f) Spearm corr = 0.98 (g) Spearm corr = 0.94 (h) Spearm corr = 0.98
Figure 7: Comparison between prevalence of synthetic data and prevalence of real dataset $\mathcal{D}_{2}$. It measures accuracy of the marginal distribution for each ICD code on synthetic samples. The second row represents the prevalence of the first row in low data regime. The prevalence is computed on 200K samples. The dash diagonal lines represent the perfect matching of code prevalence between synthetic data and real EHR data. The Pearson correlations for four methods are all greater than 0.99 and will not be used as a metric to evaluate different methods.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Density comparison of per-record feature number between synthetic data and real dataset $\mathcal{D}_{2}$. The number of feature per record is computed by sum of ICD codes on each sample. Number of bins is 90 and the range of feature number values is (0, 90).
Table 5: Synthetic data utility. Disease prediction from ICD codes on real dataset $\mathcal{D}_{2}$. AUPRC and AUROC are reported. AUPR and AUC in table are short for AUPRC and AUROC respectively. We use synthetic data of 160K to train the classifier and 200K real test data to evaluate different methods. 80% of test data are bootstrapped 50 times to compute for 95% confidence interval (CI). The values of CI for all cases are between 0.001 and 0.011.
T2D Asthma COPD CKD HTN-Heart Osteoarthritis
AUPR AUC AUPR AUC AUPR AUC AUPR AUC AUPR AUC AUPR AUC
Real data 0.834  0.955 0.581  0.853 0.622  0.951 0.733  0.944 0.278  0.926 0.373  0.893
Med-WGAN 0.725  0.924 0.496  0.819 0.203  0.853 0.166  0.835 0.008  0.500 0.223  0.820
EMR-WGAN 0.734  0.918 0.431  0.747 0.402  0.888 0.628  0.907 0.134  0.844 0.210  0.753
EHRDiff 0.807  0.950 0.549  0.843 0.548  0.936 0.690  0.916 0.141  0.822 0.319  0.875
EHR-D3PM 0.821  0.952 0.572  0.853 0.607  0.947 0.714  0.944 0.226  0.911 0.348  0.889
@@ B.2 Additional experiments on guided generation
We provide additional experiment results on guided generation. We augment the real data with synthetic data generated by our sampling method and train a downstream classifier. We measure the performance of all classifiers on real test data. From Fig. 9 and Fig. 5, we see that the classifier trained on augmented training data with synthetic data, either by our unconditional sampling method or by our guided sampling method, consistently outperforms the classifier trained with original source data (vanilla baseline). In all cases, the relative increase of AUPRC over vanilla baseline is more than 3%; in the classification of COPD, the relative improvement over the vanilla baseline is more than 30%. This clearly indicates that our method can be applied to augment the training data of downstream classification tasks when the real dataset is scarce. More importantly, the data augmentation by guided sampler consistently outperforms the data augmentation with unconditional sampler. We observe this to be because the synthetic samples generated by guided sampling contain richer information in diseases of low prevalence. A most balanced training data with positive label will enhance the performance of classifiers and reduce the risk of over fitting to negative class.
@@
T2D Asthma COPD CKD HTN-Heart Osteoarthritis
Med-WGAN 0.924 $\pm$ 0.001 0.819 $\pm$ 0.002 0.853 $\pm$ 0.003 0.835 $\pm$ 0.004 0.5 $\pm$ 0.001 0.820 $\pm$ 0.003
Table 10: Fidelity metrics (CMD and MMD) on MIMIC, dataset $\mathcal{D}_{1}$ and dataset $\mathcal{D}_{2}$. 95% confidence intervals are provides.
CMD($\downarrow$) MMD($\downarrow$)
MIMIC              $\mathcal{D}_{1}$               $\mathcal{D}_{2}$ MIMIC             $\mathcal{D}_{1}$              $\mathcal{D}_{2}$
Med-WGAN 27.540$\pm$0.628  18.107$\pm$0.128  28.942$\pm$0.196 0.078$\pm$0.0089 0.075$\pm$0.011 0.086$\pm$0.013
EMR-WGAN 26.658$\pm$0.638  11.869$\pm$0.108  21.438$\pm$0.146 0.053$\pm$0.0054 0.018$\pm$0.003 0.024$\pm$0.004
EHRDiff 25.447$\pm$0.488  23.208$\pm$0.088 18.941$\pm$0.092 0.009$\pm$0.0013 0.023$\pm$0.003 0.046$\pm$0.005
EHR-D3PM 21.128$\pm$0.393  7.692$\pm$0.028  10.255$\pm$0.037 0.003$\pm$0.0004 0.012$\pm$0.001 0.019$\pm$0.002
Table 11: Additional fidelity metric (MCAD) on MIMIC, dataset $\mathcal{D}_{1}$ and dataset $\mathcal{D}_{2}$.
MCAD($\downarrow$)
MIMIC                     $\mathcal{D}_{1}$                      $\mathcal{D}_{2}$
Med-WGAN 0.1896$\pm$0.0024       0.1871$\pm$0.0016      0.1944$\pm$0.0017
EMR-WGAN 0.1546$\pm$0.0167       0.1625$\pm$0.0013      0.1572$\pm$0.0015
EHRDiff 0.1439$\pm$0.0015      0.1687$\pm$0.0014      0.1764$\pm$0.0016
EHR-D3PM 0.1013$\pm$0.0010      0.0873$\pm$0.0007      0.1081$\pm$0.0009
Table 12: Privacy metric (MIR) on MIMIC, dataset $\mathcal{D}_{1}$ and dataset $\mathcal{D}_{2}$.
Privacy MIR($\downarrow$)
MIMIC       $\mathcal{D}_{1}$          $\mathcal{D}_{2}$
Med-WGAN 0.440$\pm$0.0034      0.339$\pm$0.0018      0.398$\pm$0.0019
EMR-WGAN 0.456$\pm$0.0035      0.358$\pm$0.0017      0.415$\pm$0.0020
EHRDiff 0.445$\pm$0.0034      0.353$\pm$0.0015      0.421$\pm$0.0019
EHR-D3PM 0.432$\pm$0.0034      0.344$\pm$0.0016      0.406$\pm$0.0018

### v2

@@ 6 Conclusion and Future Work
In this paper, we introduced a novel generative model for synthesizing realistic EHRs EHR-D3PM. Leveraging the latest advancements in discrete diffusion models, EHR-D3PM overcomes the challenges of GAN-based approaches and effectively generates high-quality tabular medical data. Compared with other diffusion-based approaches, EHR-D3PM enables high-quality conditional generation. Our experiment demonstrates that EHR-D3PM not only achieves state-of-the-art performance in terms of fidelity, utility, and privacy metrics but also significantly improves downstream task performance through data augmentation. Incorporating longitudinal feature into our model is an interesting research direction. Further investigations of the vulnerability of diffusion-based generative models in EHR generation, particularly to attribute and membership inference attacks (Shokri et al., 2017), is a promising future direction as well as providing formal privacy guarantees, e.g., by incorporating differential privacy, which is largely unexplored in diffusion-based models for discrete data.
Broader Impacts
The primary goal of this work was to develop a state-of-the-art generative model for medical code data in EHRs. The synthetic data generated by this model may be used to train, and augment the training of various downstream models for computational medicine. We believe this work may enable future applications of machine learning to healthcare which lead to positive impacts for patients, e.g., through early prediction of disease progression or matching patients to appropriate clinical trials that advance medicine. While the use of generative models and synthetic data can enhance patient privacy, they do not eliminate the risks of membership inference and other attacks, as we mention and demonstrate in the paper. Our experiments demonstrate that the proposed model achieves state-of-the-art performance in terms of reducing these risks, but proper controls must still be used whenever working with patient data. Furthermore, any downstream models trained using synthetic data must go through the same rigorous review and testing process to prevent disparate impact as models trained on real data.
Limitations
Our proposed EHR-D3PM model focuses on generating synthetic tabular EHRs and achieves state-of-the-art performance, which has a significant impact on various tasks. However, there are certain limitations to our work that should be acknowledged.
Firstly, our model does not currently incorporate longitudinal features, which could potentially enhance the utility of the generated data. Incorporating temporal dependencies and modeling the evolution of patient records over time is an important direction for future research. This extension would enable the generation of more realistic and comprehensive EHR data, capturing the dynamic nature of patient health over extended periods.
Secondly, the vulnerability of diffusion-based generative models in EHR generation to more advanced privacy attacks is a largely unexplored area that requires further research. While our model demonstrates strong privacy preservation properties, it is crucial to investigate its resilience against sophisticated adversarial attacks specifically designed to target EHR data. Thorough analysis and development of robust defense mechanisms are necessary to ensure the long-term security and privacy of the generated synthetic EHRs.
By acknowledging these limitations, we aim to provide a transparent assessment of our work’s scope and potential areas for future exploration.
@@
In the following Table 5, we present a concise summary of various diseases along with their corresponding International Classification of Diseases, Ninth Revision (ICD 9) codes. This table includes common conditions such as Type II Diabetes (T2D), Chronic Kidney Disease (CKD), Chronic Obstructive Pulmonary Disease (COPD), Asthma, Hypertension and Osteoarthritis. Each disease is associated with specific ICD 9 codes that are used for clinical classification and diagnosis purposes. In this paper, we are interested in the diseases listed in Table 5.
@@
Disease ICD 9 Code MIMIC $\mathcal{D}_{1}$ $\mathcal{D}_{2}$
@@ A.1 Dataset Details
The MIMIC III dataset includes a patient population of 46,520. There are 651,047 positive codes within 64,314 hospital admission records (HADM IDs). We have implemented an 80/20 split for training and testing purposes. Specifically, this allocates 12,862 records for testing and the remaining 51,451 for training. The histograms in Figure 6 indicate the density distribution of feature number per record. The dimension is $N=1042.$
@@
The first dataset, denoted by $\mathcal{D}_{1}$, includes a patient population of size 1,670,347. We split the whole dataset into 100K for validation, 2000K for testing and the rest 1,370, 347 for training. The number of codes per patient is relatively small, as indicated by the histogram of feature number per record in Figure 8. We only consider medical codes with prevalence in corpus larger than 1.6e-5. The dimension is $N=993.$
@@
The second dataset, denoted by $\mathcal{D}_{2}$, includes a patient population of size 1,859,536. We split the whole dataset into 100K for validation, 2000K for testing and the rest 1,559,536 for training. The number of codes per patient is relatively large, as indicated by the histogram of feature number per record in Figure 2. Although, dataset $\mathcal{D}_{2}$ has relatively denser feature, the prevalence of six chronic diseases we are interested in is pretty low. $\mathcal{D}_{2}$ contains diagnose codes (ICD), procedure codes (CPT) and medication codes (GEN). The number of ICD codes, CPT codes and GEN codes are 993, 690 and 1000 respectively. The total dimension is 2683.
@@ A.2 Model Architecture Detail
The denoise model in this paper has a uniform architecture, illustrated in Figure 4. The architecture we propose is tailored for tabular EHRs as it is non-sequential data. While the architecture proposed in multinomial diffusion Hoogeboom et al. (2021b) is designed for sequential data, where neighboring dimensions have semantic correlation. The tabular EHR datasets in our paper don’t have such property. Therefore, we propose a novel transformer-based model for tabular EHRs. One bottleneck of transformer models is that the computational complexity of the attention module is quadratic to the dimension of input data. We adopt an efficient block based on Wang et al. (2020), whose attention operation has linear complexity with respect to the dimension of input data.
@@ A.3 Hyper-parameters
In the optimization phase, we adopt adamW optimizer, and the weight decay in adamW is 1.e-5. The learning rate is 1e-4 and batch size is 256. The beta for expentialLR in learning rate schedule is 0.99. The number of training epochs is 100. It takes less than three hours to finish training this model on A6000 with 48G memory.
@@
The optimization parameters for both datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are also the same. In the optimization phrase, we adopt adamW optimizer. The learning rate is 1.0e-4 and batch size is 512. The weight decay in adamW is 1.0e-5. The beta for expentialLR in learning rate schdule is 0.99. The number of training epochs is 40. It takes one and half day to train one model on A100 with 80G memory.
@@
Hyper-parameters of baselines TabSyn and TabDDPM
To have a fair comparison with baselines TabSyn and TabDDPM, the number of diffusion steps in TabSyn and TabDDPM are also 500 and the number of layers in denoising models of TabSyn and TabDDPM is also 5. For other hyper parameters, we use the default parameters provided on the official implementation of TabSyn and TabDDPM. When training TabSyn on our datasets, we found that TabSyn works quite well in the first 100 dimensions of our datasets. But training TabSyn in the first 200 dimensions of our datasets, we found the performance significantly degenerates. The dimensions of seven datasets used in the paper of TabSyn are all less than 50.
@@ A.4 Evaluation Metrics
$\widehat{\mathrm{MMD}}_{k_{\gamma}}(P,Q)=\frac{1}{n(n-1)}\bigg{[}\sum_{i\neq j}k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})+\sum_{i\neq j}k_{\gamma}(\mathbf{x}^{\prime}_{i},\mathbf{x}^{\prime}_{j})\bigg{]}-\frac{1}{n^{2}}\sum_{i,j}k_{\gamma}(\mathbf{x}_{i},\mathbf{x}^{\prime}_{j}).$
@@
$k_{\gamma}(\mathbf{x},\mathbf{x}^{\prime})=\mathrm{exp}\bigg{(}-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2h^{2}_{\gamma}}\bigg{)}$
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Spearman corr = 0.54 (b) Spearman corr = 0.89 (c) Spearman corr = 0.75 (d) Spearman corr = 0.94
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Spearman corr = 0.45 (f) Spearman corr = 0.88 (g) Spearman corr = 0.71 (h) Spearman corr = 0.93
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Density comparison of per-record feature number on synthetic and real data for the MIMIC dataset. The number of features per record is the sum of ICD codes present in each sample. The number of bins is 40, and the range of feature number values is (0, 40).
@@ B.1 Additional experiments on unconditional generation
Figure 5 and Figure 7 provide additional comparison of marginal distribution matching on dataset MIMIC and dataset $\mathcal{D}_{1}$. We have found that TabSyn works well in low dimension but has poor performance on our high dimensional datasets with sparse features. From the Spearman correlation in low prevalence regime, we can still see that our method significantly outperform baselines. Based on results in Figure 5, Figure 7 and Figure 1, we consistently observe synthetic data by EHRDiff fails to capture the information in low prevalence regime. One reason we articulate is that the foundation of EHRDiff is designed for continuous distribution and cannot be readily applied to the generation of discrete data particularly on data of low prevalence regime. From Figure 6 and Figure 8, we can see that the histogram of feature number per record on synthetic data by our method provides the best matching to that of real data.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Spearman corr = 0.69 (b) Spearman corr = 0.97 (c) Spearman corr = 0.85 (d) Spearman corr = 0.99
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Spearman corr = 0.61 (f) Spearman corr = 0.95 (g) Spearman corr = 0.79 (h) Spearman corr = 0.97
Refer to caption Refer to caption Refer to caption Refer to caption
@@
We also apply our methods to downstream prediction tasks on dataset $\mathcal{D}_{1}$, where the prevalence of six chronic diseases is much lower. From Table  6 and Table 7, we can see that the accuracy of our prediction is still close to the prediction of classifier models trained on real data, which is the target classifier baseline. While other baselines have a much larger performance gap when compared with the ideal classifier. Particularly on rare diseases such as hypertension heart, the classifier trained on synthetic data by our model has 8% absolute improvement in AUPRC and AUROC over the strongest baseline on both $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. From the confidence intervals provided in Table 6, 7, 2 and 8, we confirm that such improvement over the baselines on both dataset $\mathcal{D}_{1}$ and dataset $\mathcal{D}_{2}$ are statistically significant.
@@ B.2 Additional experiments on guided generation
We provide additional experiment results on guided generation. We augment the real data with synthetic data generated by our sampling method and train a downstream classifier. We measure the performance of all classifiers on real test data. From Figure 9 and Figure 3, we see that the classifier trained on augmented training data with synthetic data, either by our unconditional sampling method or by our guided sampling method, consistently outperforms the classifier trained with original source data (vanilla baseline). In all cases, the relative increase of AUPRC over vanilla baseline is more than 3%; in the classification of COPD, the relative improvement over the vanilla baseline is more than 30%. This clearly indicates that our method can be applied to augment the training data of downstream classification tasks when the real dataset is scarce. More importantly, the data augmentation by guided sampler consistently outperforms the data augmentation with unconditional sampler. We observe this to be because the synthetic samples generated by guided sampling contain richer information in diseases of low prevalence. A most balanced training data with positive label will enhance the performance of classifiers and reduce the risk of over fitting to negative class.
@@
Table 9: Fidelity metrics (CMD, MMD and MCAD) on dataset $\mathcal{D}_{1}$.
Metrics CMD($\downarrow$) MMD($\downarrow$) MCAD($\downarrow$)
Med-WGAN 18.107$\pm$0.125 0.075$\pm$0.011 0.1871$\pm$0.0016
EMR-WGAN 11.869$\pm$0.108 0.018$\pm$0.003 0.1625$\pm$0.0013
TabDDPM 11.204$\pm$0.164 0.021$\pm$0.005 0.1542$\pm$0.0019
EHRDiff 23.208$\pm$0.083 0.023$\pm$0.003 0.1687$\pm$0.0014
EHR-D3PM 7.692$\pm$0.026 0.012$\pm$0.001 0.0873$\pm$0.0007